A table tennis ball has a diameter of 3.80 cm and an average density of 0.0840 g/cm3. What force is required to hold it completely submerged underwater?


1
Expert's answer
2022-03-20T18:58:01-0400

Given

"d-" Diameter of a tenis ball=3.80cm

"\rho-" Density of the tenis ball"=0.0840g\/cm^3"

"dw-" Density of water"=1g\/cm^3"

Volume of the ball,"w-v=43 \pi d38"

"=16\pi\times(3.8)3cm^3=28.73cm^3"

"w-" weight of the ball"=v×\rho ×g"

when the ball is submerged the buoyant force acting on the ball wil be equal to the weight of displaced water.

So the bouyant force

"Fb=v\times dw \times g"

So the net downward force to be appliied on the ball to just submerge it will be

"F_{net}=Fb-w=vdwg-v\rho g=vg(dw-\rho )"

"=28.73 \times 980 \times (1-0.084) dyne = 25790.3 dyne"